# From Uncertainty to Trust: Enhancing Reliability in Vision-Language Models with Uncertainty-Guided Dropout Decoding

### v1

@@ 5.1 Uncertainty Quantification Before Decoding
where $q^{\text{proj}}_{i}$ represents the text-space projection of the $i$-th visual token, and $N$ is the total number of visual tokens. Note that the subscript $i$ indicates different distributions rather than elements within a single distribution. This provides us with a “baseline” representation of the visual input, against which we can quantify the surprisal of a specific visual token. This idea is grounded in classical uncertainty decomposition where a Bayesian average distribution is needed to quantify epistemic uncertainty [20, 43].
Uncertainty measurement for visual tokens. We aim to quantify the uncertainty associated with each visual token at inference time. To distinguish from those uncertainty terms in classical settings as introduced in §3.2, we refer to ours as perception uncertainty. We start by quantifying the perception total uncertainty of the visual input as the entropy of the average visual-textual distribution $\mathbf{\mathbb{H}}\left[q^{\text{proj}}\right]$. Then, to attribute this total uncertainty to individual visual tokens, we decompose it (more details in Appendix A) as follows:
@@ 5.2 Uncertainty-Guided Decoding
Identifying relevant visual tokens (optional). We selectively retain only the most relevant visual tokens from the context, which are excluded for dropout. To do this, when generating each output text token, $y_{j}$, we first perform a preliminary forward pass to generate an initial prediction token $y^{\text{init}}_{j}$:
@@
It is worth noting that this step is optional. Omitting it can improve efficiency by reducing the computational overhead of the preliminary forward pass. As shown by the ablation studies in §7, while skipping this step may lower performance on certain benchmarks like THRONE [23], it still achieves comparable results on others such as CHAIR [41].
Visual token dropout with uncertainty guidance. Using the epistemic uncertainty measurements $U_{\text{epi}}(i)$ from Eq. (6), we introduce dropout masks over visual tokens. As illustrated in Fig. 1, the projected visual-textual distributions sometimes misalign with the image content, and regions of high information can lead to substantial errors, resulting in hallucinations. Building on this intuition, we selectively target visual tokens with high epistemic uncertainties for dropout.
@@
where $U_{\text{epi}}^{\min}$, $U_{\text{epi}}^{\max}$ are the minimum and maximum epistemic uncertainty values across all visual tokens, and $\gamma^{(k)}$ and $\delta^{(k)}$ are hyperparameters controlling the probability range of the dropout. By adjusting the values of $\gamma^{(k)}$ and $\delta^{(k)}$, we can modulate the intensity of visual token dropout.
With the dropout distributions, we can sample dropout masks for each visual token independently. Denote the binary mask as $M^{(k)}\in\{0,1\}^{N}$, consisting of a binary indicator $M^{(k)}_{i}$ for each visual token $x^{v}_{i}$, where the conrresponding visual token is retained if $M^{(k)}_{i}=1$, and dropped if $M^{(k)}_{i}=0$. The dropout mask sampling follows $P(M^{(k)}_{i}=0)=P^{(k)}_{\text{dropout}}(x^{v}_{i})$, and the sampling is done for each visual token position independently. A higher value of $P_{\text{dropout}}(x^{v}_{i})$ indicates that $x^{v}_{i}$ is more likely to be dropped out. If we performed the optional preliminary forward pass to identify relevant visual token set $\mathcal{S}_{j}$, these visual tokens are never dropped, i.e., $\forall x^{v}_{i}\in\mathcal{S}_{j}$, set $M^{(k)}_{i}=1$ directly.
@@
[…]
Each $y_{j}^{(k)}$ serves as a candidate prediction for the next text token, with the final token $y_{j}$ selected via majority voting among the $K$ masked inputs. In case of a tie, we choose the prediction from the forward pass with the fewest dropped tokens, as it retains the most information and is deemed more reliable. By forming an ensemble of predictions derived from various subsets of the visual input, enabled through token dropout, we diversify the model’s perspective on the visual content. This diversity mitigates the impact of any single misinterpretation, ultimately leading to more reliable and robust generation, which is also observed in other ensemble-based methods [4, 42, 14, 25, 12].
@@
5:  Obtain visual text projecting distributions $q^{\text{proj}}_{i}$. $\triangleright$ Eq (3)
6:  Compute average distribution $q^{\text{proj}}$. $\triangleright$ Eq. (4)
7:  Compute epistemic uncertainty $U_{\text{epi}}(i)$. $\triangleright$ Eq. (6)
@@
10:     Generate preliminary token $y^{\text{init}}_{j}$. $\triangleright$ Eq. (7)
11:     Get relevant tokens $\mathcal{S}_{j}$ with $y^{\text{init}}_{j}$ and $q^{\text{proj}}_{i}$. $\triangleright$ Eq. (8)
12:
13:     Visual token dropout with uncertainty-guidance:
14:     Get $K$ dropout prob $P^{(k)}$ with $U_{\text{epi}}(i)$. $\triangleright$ Eq. (9)
15:     Generate $K$ dropout masks $M^{(k)}$ based on $P^{(k)}$ while retain relevant tokens $\mathcal{S}_{j}$.
16:     Forward candidates $y^{(k)}_{j}$ with masks $M^{(k)}$. $\triangleright$ Eq. (10)
17:     Majority voting on $y^{(k)}_{j}$ and get $y_{j}$.
18:  end for
19:  Return Generated sequence $y$
@@ 6 Experiments
Dropout Decoding 26.26±2.4 7.39±0.69 0.821±0.010 0.840±0.009 0.842±0.002 0.805±0.010
Dropout Decoding (w/o prelim) 27.0±1,80 7.53±0.643 0.814±0.009 0.835±0.007 0.837±0.003 0.793±0.008
Table 1: Comparison of methods on CHAIRS, CHAIRI, $F^{1}_{\text{all}}$, $F^{0.5}_{\text{all}}$, $P_{\text{all}}$, and $R_{\text{all}}$ metrics for LLaVA-1.5, InstructBLIP, and LLaVA-NEXT.
We evaluate the proposed Dropout Decoding from two aspects: OH reduction and overall generation quality. For OH, we use the CHAIR [41] and THRONE [23] metrics to assess the performance of different decoding methods on the MSCOCO dataset. Additionally, we employ MMBench [36] to evaluate the overall generation quality and general ability of these methods.
6.1 Experimental Setup
Base LVLMs. We evaluate all methods on three representative LVLMs: LLaVA-1.5 [32], InstructBLIP [9] and LLaVA-NEXT  [34]. LLaVA-1.5 employs linear projection layers to align image and text features, generating 576 visual tokens for detailed visual representation, while LLaVA-NEXT extends this approach by utilizing thousands of visual tokens. In contrast, InstructBLIP uses a Q-former with only 32 visual tokens to bridge the modalities. This diversity highlights the flexibility of our approach, validating its effectiveness across models with both high and low token counts, and confirming its robustness and adaptability.
Hallucination reduction baselines. In addition to the original LVLM outputs, we compare our method with beam search as well as two state-of-the-art decoding methods: VCD [27], which contrasts original and distorted visuals to reduce hallucinations, and OPERA [19], which applies penalties and token adjustments for better grounding.
@@ 6.2 CHAIR
CHAIR [41] is a benchmark designed to evaluate OH in image captioning situations. CHAIR provides two primary metrics to measure hallucination at different granularities: sentence-level and object-level. The sentence-level metric, $\text{CHAIR}_{S}$, calculates the proportion of captions that contain any hallucinated objects, giving an overall measure of hallucination frequency in captions. And the object-level metric, $\text{CHAIR}_{I}$, calculates the proportion of hallucinated objects out of all mentioned objects across captions, reflecting the prevalence of hallucination among the objects described.
Results. As shown in Table 1, Dropout Decoding consistently outperforms baseline approaches across various models, demonstrating its reliability and effectiveness in image captioning. Especially, on InstructBLIP, CHAIRI is improved by approximately  16% over the second-best method, and CHAIRS sees a gain of around  12%. These substantial improvements underscore the effectiveness of our approach, which aligns well with intuitive expectations that token dropout will reduce generated objects. Furthermore, Dropout Decoding reduces the generation of hallucinated objects without compromising the inclusion of relevant objects. This reduction in hallucinated content, as opposed to accurate content, is further validated by the recall metric ($R_{\text{all}}$) in THRONE.
@@ 6.3 THRONE
THRONE [23] assesses hallucinations in LVLM-generated responses, covering both “Type I” (mentions of non-existent objects, like CHAIR) and “Type II” (accuracy of object existence, like POPE [30]). It uses $P_{\text{all}}$ (Precision), $R_{\text{all}}$ (Recall), $F^{1}_{\text{all}}$, and $F^{0.5}_{\text{all}}$. Additionally, it employs Fβ, which combines $P_{\text{all}}$ and $R_{\text{all}}$, with the parameter $\beta$ controlling the weight of $R_{\text{all}}$ relative to $P_{\text{all}}$: $F^{\beta}_{\text{all}}=(1+\beta^{2})\cdot\frac{P_{\text{all}}\times R_{\text{all}}}{(\beta^{2}\times P_{\text{all}})+R_{\text{all}}}.$
@@ 6.4 MMBench
MMBench [36] is a comprehensive benchmark designed to evaluate the multimodal capabilities of LVLMs across various tasks and data types, including image captioning, question answering, and object recognition. It provides a holistic view of a model’s strengths and weaknesses in multimodal understanding. Since the prompt length limits in MMBench exceed InstructBLIP’s token allowance, we report results only on LLaVA-1.5 and LLaVA-NEXT.
@@
Table 2: Results of different methods on MMBench.
@@
Results. As shown in Table 2, Dropout Decoding outperforms all the other baselines on LLaVA-1.5, which demonstrates not only its effectiveness in hallucination mitigation but also its robustness and adaptability across a broader range of multimodal tasks.
@@ 7 Analysis and Ablation Studies
7.1 Number of Parallel Dropouts
As described in §5.2, we generate $K$ candidate predictions by applying token dropout with different dropout masks $M^{(k)}$. In this section, we investigate how varying $K$ from 1 to 4 affects generation quality.
@@ 7.1 Number of Parallel Dropouts
We fix $\delta^{(k)}=0.1$ and adjust $\gamma^{(k)}$ based on a predefined order: $\gamma^{(1)}=0.3$, $\gamma^{(2)}=0.5$, and $\gamma^{(3)}=0.7$. However, setting $\gamma^{(k)}$ to 0.9 leads to excessive dropout of visual tokens and degrades InstructBLIP’s performance, so we set $\gamma^{(4)}=0.1$. Moreover, our majority voting algorithm favors candidates with fewer dropped tokens in the event of a tie, meaning that when comparing only two candidates, both will yield identical outputs. To address this, we remove Candidate 1 in the second round, leaving only Candidate 2.
@@
As shown in Fig. 3 (a) and (b), both CHAIRS and CHAIRI scores peak at $K=3$ for LLaVA-1.5 and InstructBLIP. Increasing $K$ to 4 introduces a less-masked candidate that slightly negatively impact our method’s effectiveness in reducing hallucinations. Conversely, using fewer candidates (e.g., only Candidate 1 and Candidate 2) lacks the balance needed for stable voting outcomes, resulting in increased randomness. Similarly, Fig. 3 (c) and (d) shows that THRONE’s $R_{\text{all}}$ and $P_{\text{all}}$ metrics also perform best at $K=3$. Overall, we find that selecting three candidates strikes the optimal balance between increased certainty from additional votes and the controlled uncertainty introduced by candidate dropout probability, allowing Dropout Decoding to achieve more trustworthy and stable generation results.
7.2 Initial Identification of Relevant Visual Tokens
@@
As discussed in §5.2, Dropout Decoding may employ a preliminary forward pass to retain most relevant objects during generation, which helps reduce hallucinated objects while maintaining high-quality outputs. In contrast, bypassing this step risks masking relevant visual tokens during the token dropout phase, potentially degrading overall performance. However, incorporating a preliminary forward pass roughly doubles the computational cost per generation. To strike a balance between accuracy and efficiency, we evaluate the approach both with and without this step. Specifically, our goals are: 1) to confirm the effectiveness of the preliminary forward pass, and 2) to explore a more efficient alternative when computational resources are limited.
@@
As shown in Table 1, including the preliminary forward pass consistently improves most metrics, with particular notable gains in the $F_{\text{all}}$ score on THRONE. We also observe further improvements in CHAIR metrics, especially when using InstructBLIP. Interestingly, for LLaVA-1.5, the variant without the preliminary pass performs slightly better on CHAIR, though the THRONE performance remains stable (i.e., the version with the preliminary pass still performs better). We hypothesize that this discrepancy arises from differences in the abundance of visual tokens. LLaVA-1.5 uses 576 visual tokens, each carrying relatively less weight, whereas InstructBLIP relies on just 32, making each token’s contribution more critical. Consequently, omitting the preliminary forward pass in InstructBLIP risks losing critical information, lowering performance. These findings suggest that while a preliminary forward pass is highly beneficial for models with abundant visual tokens, models with fewer tokens may achieve better computational efficiency and performance by skipping this additional step.

### v2

@@ 5.1 Uncertainty Quantification Before Decoding
where $q^{\text{proj}}_{i}$ represents the text-space projection of the $i$-th visual token, and $N$ is the total number of visual tokens. Note that the subscript $i$ indicates different distributions rather than elements within a single distribution. This provides us with a “baseline” representation of the visual input, against which we can quantify the surprisal of a specific visual token. This idea is grounded in classical uncertainty decomposition where a Bayesian average distribution is needed to quantify epistemic uncertainty (hullermeier2021aleatoric; schweighofer2023introducingimprovedinformationtheoreticmeasure).
Uncertainty measurement for visual tokens. We aim to quantify the uncertainty associated with each visual token at inference time. To distinguish from those uncertainty terms in classical settings as introduced in §3.2, we refer to ours as perception uncertainty. We start by quantifying the perception total uncertainty of the visual input as the entropy of the average visual-textual distribution $\mathbf{\mathbb{H}}\left[q^{\text{proj}}\right]$. Then, to attribute this total uncertainty to individual visual tokens, we decompose it (details in Appendix A) by:
@@ 5.2 Uncertainty-Guided Decoding
Identifying relevant visual tokens (optional). We selectively retain only the most relevant visual tokens from the context, which are excluded for dropout. When generating each output text token, $y_{j}$, we first perform a preliminary forward pass to generate an initial prediction token $y^{\text{init}}_{j}$:
@@
It is worth noting that this step is optional. Omitting it can improve efficiency by reducing the computational overhead of the preliminary forward pass. As shown by the ablation studies in §7, while skipping this step may lower performance on certain benchmarks like THRONE (kaul2024throneobjectbasedhallucinationbenchmark), it still achieves comparable results on others such as CHAIR (rohrbach2019objecthallucinationimagecaptioning).
Visual token dropout with uncertainty guidance. Using the epistemic uncertainty measurements $U_{\text{epi}}(i)$ from Eq. (6), we introduce dropout masks over visual tokens. As illustrated in Fig. 1, the projected visual-textual distributions sometimes misalign with the image content, and regions of high information can lead to substantial errors, resulting in hallucinations. Based on this intuition, we selectively target visual tokens with high epistemic uncertainties for dropout.
@@
where $U_{\text{epi}}^{\min}$, $U_{\text{epi}}^{\max}$ are the minimum and maximum epistemic uncertainty values across all visual tokens, and $\gamma^{(k)}$ and $\delta^{(k)}$ are hyperparameters controlling the probability range of the dropout. By adjusting the values of $\gamma^{(k)}$ and $\delta^{(k)}$, we can modulate the intensity of visual token dropout. For further discussion on hyperparameters, see §7.2.
With the dropout distributions, we can sample dropout masks for each visual token independently. Denote the binary mask as $M^{(k)}\in\{0,1\}^{N}$, consisting of a binary indicator $M^{(k)}_{i}$ for each visual token $x^{v}_{i}$, where the conrresponding visual token is retained if $M^{(k)}_{i}=1$, and dropped if $M^{(k)}_{i}=0$. The dropout mask sampling follows $P(M^{(k)}_{i}=0)=P^{(k)}_{\text{dropout}}(x^{v}_{i})$, and the sampling is done for each visual token position independently. A higher value of $P_{\text{dropout}}(x^{v}_{i})$ indicates that $x^{v}_{i}$ is more likely to be dropped out. If we performed the optional preliminary forward pass to identify relevant visual token set $\mathcal{S}_{j}$, these visual tokens are never dropped, i.e.,, $\forall x^{v}_{i}\in\mathcal{S}_{j}$, set $M^{(k)}_{i}=1$ directly.
@@
$y^{(k)}_{j}\overset{\text{Decoding}}{\sim}p_{\theta}(\cdot\mid x^{v}_{/M^{(k)}},x^{t},y_{<j})$ (10)
where $x^{v}_{/M^{(k)}}$ denotes the visual tokens after applying dropout mask $M^{(k)}$, and $\overset{\text{Decoding}}{\sim}$ denotes invariance to the decoding algorithm used (e.g., greedy search in our implementation, though others are applicable).
Each $y_{j}^{(k)}$ serves as a candidate prediction for the next text token, with the final token $y_{j}$ selected via majority voting among the $K$ masked inputs. In case of a tie, we choose the prediction from the forward pass with the fewest dropped tokens, as it retains the most information and is deemed more reliable. By forming an ensemble of predictions derived from various subsets of the visual input, enabled through token dropout, we diversify the model’s perspective on the visual content. This diversity mitigates the impact of any single misinterpretation, ultimately leading to more reliable and robust generation, which is also observed in other ensemble-based methods (chen2024halcobjecthallucinationreduction; rokach2010ensemble_based_classifiers; ganaie2022ensemble_deep_learning_review; lakshminarayanan2017simple_and_scalable_predictive_uncertainty_estimation_deep_ensembles; fort2019deep_ensembles_loss_landscape_perspective).
@@
5:  Obtain visual text projecting distributions $q^{\text{proj}}_{i}$. {Eq (3)}
6:  Compute average distribution $q^{\text{proj}}$. {Eq. (4)}
7:  Compute epistemic uncertainty $U_{\text{epi}}(i)$. {Eq. (6)}
@@
10:   Generate preliminary token $y^{\text{init}}_{j}$. {Eq. (7)}
11:   Get relevant tokens $\mathcal{S}_{j}$ with $y^{\text{init}}_{j}$ and $q^{\text{proj}}_{i}$. {Eq. (8)}
12:   Visual token dropout with uncertainty-guidance:
13:   Get $K$ dropout prob $P^{(k)}$ with $U_{\text{epi}}(i)$. {Eq. (9)}
14:   Generate $K$ dropout masks $M^{(k)}$ based on $P^{(k)}$ while retain relevant tokens $\mathcal{S}_{j}$.
15:   Forward candidates $y^{(k)}_{j}$ with masks $M^{(k)}$. {Eq. (10)}
16:   Majority voting on $y^{(k)}_{j}$ and get $y_{j}$.
17:  end for
18:  Return Generated sequence $y$
@@ 6 Experiments
We evaluate the proposed Dropout Decoding from two aspects: OH reduction and overall generation quality. For OH, we use the CHAIR (rohrbach2019objecthallucinationimagecaptioning) and THRONE (kaul2024throneobjectbasedhallucinationbenchmark) metrics to assess the performance of different decoding methods on the MSCOCO dataset. Additionally, we employ MMBench (liu2024mmbenchmultimodalmodelallaround) to evaluate the overall generation quality and general ability of these methods.
6.1 Experimental Setup
Base LVLMs. We evaluate all methods on three representative LVLMs: LLaVA-1.5 (liu2023llava), InstructBLIP (dai2023instructblipgeneralpurposevisionlanguagemodels) and LLaVA-NEXT  (liu2024llavanext). LLaVA-1.5 and LLaVA-NEXT use hundreds to thousands of visual tokens for detailed representation, while InstructBLIP employs just 32 tokens but with higher information density. This showcases the flexibility of our approach, effective across models with varying token counts.
@@ 6.1 Experimental Setup
[…]
Table 1: Comparison of methods on CHAIRS, CHAIRI, $F^{1}_{\text{all}}$, $F^{0.5}_{\text{all}}$, $P_{\text{all}}$, and $R_{\text{all}}$ metrics for LLaVA-1.5, InstructBLIP, and LLaVA-NEXT. Details of the experimental setup and the interpretation of the standard deviation can be found in the appendix. Details of the experimental setup and the interpretation of the standard deviation can be found in the appendix.
Hallucination reduction baselines. In addition to the original LVLM outputs, we compare our method with beam search as well as two state-of-the-art decoding methods: VCD (leng2023mitigatingobjecthallucinationslarge), which contrasts original and distorted visuals to reduce hallucinations, and OPERA (huang2024operaalleviatinghallucinationmultimodal), which applies penalties and token adjustments for better grounding.
@@ 6.2 CHAIR
CHAIR (rohrbach2019objecthallucinationimagecaptioning) is a benchmark for evaluating object hallucination in image captioning. It includes two metrics: the sentence-level $\text{CHAIR}_{S}$, measuring the frequency of captions with hallucinated objects, and the object-level $\text{CHAIR}_{I}$, calculating the proportion of hallucinated objects among all objects.
Results. As shown in Table 1, Dropout Decoding consistently outperforms baseline approaches across various models, demonstrating its reliability and effectiveness in image captioning. Especially on InstructBLIP, CHAIRI and CHAIRS improve by  16% and  12% respectively over the second-best method. Furthermore, Dropout Decoding reduces the generation of hallucinated objects without compromising the inclusion of relevant objects. These improvements align with expectations that token dropout reduces generated objects.
@@ 6.3 THRONE
THRONE (kaul2024throneobjectbasedhallucinationbenchmark) assesses hallucinations in LVLM-generated responses, covering both “Type I” (mentions of non-existent objects, like CHAIR) and “Type II” (accuracy of object existence, like POPE (li2023evaluatingobjecthallucinationlarge)). It uses $P_{\text{all}}$ (Precision), $R_{\text{all}}$ (Recall), $F^{1}_{\text{all}}$, and $F^{0.5}_{\text{all}}$. Additionally, it employs Fβ, which combines $P_{\text{all}}$ and $R_{\text{all}}$, with the parameter $\beta$ controlling the weight of $R_{\text{all}}$ relative to $P_{\text{all}}$: $F^{\beta}_{\text{all}}=(1+\beta^{2})\cdot\frac{P_{\text{all}}\times R_{\text{all}}}{(\beta^{2}\times P_{\text{all}})+R_{\text{all}}}.$
@@ 6.4 MMBench
MMBench (liu2024mmbenchmultimodalmodelallaround) is a comprehensive benchmark designed to evaluate the multimodal capabilities of LVLMs across various tasks and data types. Since the prompt length limits in MMBench exceed InstructBLIP’s token allowance, we report results only on LLaVA-1.5 and LLaVA-NEXT.
@@
Table 2: Results on MMBench. Higher is better.
@@
Results. As shown in Table 2, Dropout Decoding outperforms all the other baselines on LLaVA-1.5, which demonstrates its robustness and adaptability across a broader range of multimodal tasks.
@@ 7 Analysis and Ablation Studies
7.1 Efficiency Analysis
We conducted a thorough analysis of computational overhead, measuring throughput and wall-time to evaluate efficiency. Table 3 summarizes these results. Our method introduces additional overhead primarily in two aspects: (1) a preliminary forward pass for identifying relevant visual tokens, (2) performing $K$ parallel forward passes using varied dropout masks.
The preliminary forward pass, though beneficial, is optional. Omitting it results in only approximately 7% throughput reduction compared to greedy decoding, while still consistently improving performance metrics across benchmarks. Furthermore, the method efficiently handles the $K$ parallel passes by batching identical inputs with distinct dropout masks into a single batched operation, significantly reducing additional computational overhead.
In terms of GPU memory, we verified efficiency under realistic conditions. Using vLLM and LLaVA-1.5 on 4×A800 80GB GPUs, GPU memory usage was 38.12 GB with efficient KV caching, unchanged between greedy decoding and our method without preliminary passes. Confirmatory experiments under Huggingface Transformers similarly demonstrated minimal GPU memory increase (from 14.02 GB to 15.31 GB), indicating negligible impact on inference constraints.
The cost of computing uncertainty metrics is explicitly included in our benchmarks and remains negligible. With LLaVA-1.5 and 576 image tokens, computing uncertainty adds only 73.30 ms per input, a minor cost amortized across the batched forward passes.
Overall, our approach effectively balances efficiency and performance.
Metric Greedy Ours w/o prelim Ours w/ prelim
Throughput (tok/s) $\uparrow$ 37.0 34.1 (-7.8%) 20.1 (-45.7%)
Wall-time (per 50 tok) $\downarrow$ 1.35 1.47 (+8.9%) 2.49 (+84.4%)
Table 3: Computational overhead analysis.
7.2 Parallel Dropouts Hyperparameters
@@
As in §5.2, we generate $K$ candidate predictions using token dropout. This section examines how varying the hyperparameters impacts generation quality. We fix $\delta^{(k)}=0.1$ and adjust $\gamma^{(k)}$ based on a predefined order: $\gamma^{(1)}=0.3$, $\gamma^{(2)}=0.5$, and $\gamma^{(3)}=0.7$. However, setting $\gamma^{(4)}=0.9$ excessively drops visual tokens and degrades InstructBLIP’s performance, so we set $\gamma^{(4)}=0.1$. Moreover, our majority voting favors candidates with fewer dropped tokens in ties. To avoid identical outputs when comparing only two candidates, we remove Candidate 1 in the second round.
@@
As shown in Fig. 3 (a) and (b), both CHAIRS and CHAIRI scores peak at $K=3$ for LLaVA-1.5 and InstructBLIP. Increasing $K$ to 4 introduces a less-masked candidate that slightly negatively impact our method’s effectiveness in reducing hallucinations. Conversely, using fewer candidates (e.g. , only Candidate 1/2) lacks the balance needed for stable voting outcomes, resulting in increased randomness. Similarly, Fig. 3 (c) and (d) shows that THRONE’s $R_{\text{all}}$ and $P_{\text{all}}$ metrics also perform best at $K=3$. Overall, we find that selecting three candidates strikes the optimal balance between increased certainty from additional votes and the controlled uncertainty introduced by candidate dropout probability, allowing Dropout Decoding to achieve more trustworthy and stable generation results.
7.3 Preliminary Forward Pass
@@
As discussed in §5.2, Dropout Decoding may employ a preliminary forward pass to retain most relevant objects during generation, which helps reduce hallucinated objects while maintaining high-quality outputs. In contrast, bypassing this step risks masking relevant visual tokens during the token dropout phase, potentially degrading overall performance. However, incorporating a preliminary forward pass roughly doubles the computational cost per generation. Specifically, our goals are: 1) to confirm the effectiveness of the preliminary forward pass, and 2) to explore a more efficient alternative when computational resources are limited.
@@
As shown in Table 1, including the preliminary forward pass consistently improves most metrics, with particular notable gains in the $F_{\text{all}}$ score on THRONE. Interestingly, for LLaVA-1.5 and LLaVA-NEXT, the variant without the preliminary pass still outperforms other baselines in most metrics. We hypothesize that this discrepancy arises from differences in the abundance of visual tokens, as LLaVA-1.5 and LLaVA-NEXT have hundreds or thousands of visual tokens. While InstructBLIP only has 32, making each token’s contribution more critical. Consequently, omitting the preliminary forward pass in InstructBLIP risks losing critical information, lowering performance. These findings suggest that while a preliminary forward pass is highly beneficial for LVLMs, models with more tokens may achieve better efficiency and performance by skipping this step.
7.4 Necessity of Uncertainty Guidance on Masking
As discussed in §5.1, Dropout Decoding incorporates epistemic uncertainty in the masking process. To validate the necessity of this approach, we compare it with a random masking strategy, which replaces uncertainty with a random method. As shown in table Table 4, although random masking performs better on the CHAIR metric, it struggles with BLEU and fails to compute the THRONE metric. We find that random masking causes repetitive token generation (e.g., “apple apple apple…”), artificially inflating the CHAIR score. This happens because random masking disrupts contextual information, leading to faulty generation. In contrast, our uncertainty-based approach selectively masks uncertain tokens, preserving the context and ensuring more coherent and accurate sequences.
Model Method CHAIR BLEU THRONE
CHAIRS $\downarrow$ CHAIRI $\downarrow$ BLEU $\uparrow$ THRONE $\uparrow$
LLaVA-1.5 Greedy 42.20±2.86 7.90±0.63 11.62±0.09 Table 1
Uncertainty-guided 39.80±2.3 11.73±0.25 11.64±0.12 Table 1
Random 35.93±0.90 8.57±0.63 11.51±0.12 Error
InstructBLIP Greedy 27.87±1.32 7.90±0.63 12.70±0.54 Table 1
Uncertainty-guided 24.53±1.26 6.63±0.65 12.30±0.18 Table 1
Random 23.80±1.28 5.07±0.71 10.88±0.08 Error
Table 4: Comparison of masking strategies on CHAIRS, CHAIRI, BLEU and THRONE metrics for LLaVA-1.5, InstructBLIP.
7.5 High-Confidence Token Masking Analysis
To further investigate the robustness of our uncertainty-guided masking, we conducted an additional ablation experiment focusing on the opposite condition—masking high-confidence tokens instead of low-confidence ones.
Following the identical experimental setup as in Table 1 (our main CHAIR and THRONE evaluation), we replaced low-confidence masking with high-confidence masking. The results are summarized in Table 5.
Model CHAIRS $\downarrow$ CHAIRI $\downarrow$ $F_{\text{all}}$ $\uparrow$ $F_{0.5,\text{all}}$ $\uparrow$ $P_{\text{all}}$ $\uparrow$ $R_{\text{all}}$ $\uparrow$
LLaVA 41.62 12.00 0.798 0.784 0.774 0.858
LLaVA-NEXT 27.61 7.82 0.803 0.823 0.828 0.789
InstructBLIP 29.50 9.01 0.808 0.825 0.825 0.794
Table 5: Effect of masking high-confidence tokens on CHAIR and THRONE metrics.
As shown, the results are generally worse than masking low-confidence tokens (i.e., masking high-uncertainty tokens as in our proposed method) and remain close to the greedy decoding baseline. This suggests that dropping high-confidence tokens has limited influence on generation quality—these tokens correspond to regions where the model is already certain, typically associated with background or redundant patches. Consequently, their removal produces only minor perturbations.
In contrast, masking low-confidence tokens directly influences the model’s generative process, as these tokens are uncertain yet potentially informative (often corresponding to salient or ambiguous visual regions). Masking them introduces meaningful variability, thereby improving robustness and reducing hallucination frequency. This further validates our uncertainty-guided masking strategy as both effective and theoretically grounded.